# Aperiodic sets of three types of convex polygons

### v3

@@ 1 Introduction
A tiling (or tessellation) of the plane is a collection of sets that are called tiles, which covers a plane without gaps and overlaps, except for the boundaries of the tiles. The term “tile” refers to a topological disk, whose boundary is a simple closed curve.
A tiling has periodicity if its translation by a non-zero vector coincides with it; a tiling is periodic if it coincides with its translation by two linearly independent vectors. However, in this study, the tiling with periodicity is referred to as periodic; thus, the tiling with no periodicity is referred to as non-periodic.
In general, when considering tilings, there are only a finite number of tile shapes that can be used. These finite-shape diagrams are called prototiles, and a set of prototiles admits tilings to the plane. The tiling admitted by a set of tiles (tile set) can cover the plane infinitely. A set of prototiles is aperiodic if copies of the prototiles can be assembled into tilings of the plane such that all tilings with the prototiles are non-periodic [1, 3, 16].
The most well-known problem regarding a set of prototiles (set of tiles) that are aperiodic, that is, “aperiodic set of prototiles (aperiodic set of tiles)” (hereafter abbreviated as “ASP”), is “Is there a single aperiodic prototile (with or without a matching condition), that is, one that admits only non-periodic tilings by congruent copies?” [1]. Smith et al. presented solutions to this problem for concave diagram tiles without matching conditions [9, 10]. In [1], another major problem concerning ASP was presented as follows:
@@
Figure. 1: Tiling (d) generated by an aperiodic set of prototiles created by Ammann, which used three convex polygons with no matching condition on the edges. As shown in (c), this is a recomposition by Penrose tiles (rhombuses) and is obtained by markings of using the points $J$, $K$, and $L$ as shown in (a) and (b). These markings are completely determined by the choice of the point $J$, as we must have GL = DJ, FL = CJ =EK, and HK = BJ. For aperiodicity, the point $J$ must be chosen so that DJ, CJ, BJ, and KL are of different lengths [3, 12].
The “ASP comprising three convex polygons (= the set of three convex polygons that are aperiodic with no matching condition on the edges)” in the above problem is known for the two convex pentagons and one convex hexagon created by Ammann, as shown in Figure 1, based on the Penrose tiles111 The Penrose tiles are considered to be an ASP comprising two convex polygons (rhombuses) with matching condition on the edges [1, 3, 12]. [3, 12]. In the tiling of the two convex pentagons and one convex hexagon shown in Figure 1, the reflected (flipped) convex pentagons and hexagon were not used. The tiling is edge-to-edge222 A tiling by convex polygons is edge-to-edge if any two convex polygons in a tiling are either disjoint or share one vertex or an entire edge in common. Then other case is non-edge-to-edge [3, 11, 12]..
@@
According to Smith et al. [9], Tile$(a,b)$, which can generate only non-periodic tiling, is an “aperiodic monotile,” in which all tiles are of the same size and shape; however, the tiling generation process requires the use of reflected tiles, and both anterior and posterior side tiles must be treated as the same type. However, according to Smith et al. [10], the tiles “Spectres,” which can generate only non-periodic tiling, do not require the use of reflected tiles during the tiling generation process333 Spectres can generate only non-periodic tiling using only one side, either anterior or posterior, even if it allows the use of reflected tiles during the tiling generation process.. The tiles are referred to as “chiral aperiodic monotile.”
@@
Following this policy, the case in [9] is “ASP comprising two types of concave polygons,” and the case in [10] is “ASP comprising one type of concave diagram444 Tile(1,1)11(1,1)( 1 , 1 ) corresponding to $a=b=1$ of Tile$(a,b)$ shown in [9] is a concave polygon and a concave diagram, and Spectres in [10] are concave diagrams. From a viewpoint that does not distinguish between anterior and posterior sides, both cases of [9] and [10] are considered “ASP comprising one type of concave diagram.” Note that Tile(1,1)11(1,1)( 1 , 1 ) with 13-edges is considered an equilateral 14-edges polygon.”. Moreover, the case shown by Ammann in Figure 1 is “ASP comprising three types of convex polygons (two types of convex pentagons and one type of convex hexagon).”
@@
The two types of convex pentagons and one type of convex hexagon shown in Figure 1 can be identified as not belonging to known Type families of convex pentagonal monotiles (see Appendix A) or convex hexagonal monotiles (see Appendix B), respectively555 If Rao’s claim in [7] is formally established as correct, there will be no ASP for convex polygonal monotiles (i.e., there does not exist ASP comprising one type of convex polygon with no matching condition on the edges), and furthermore, it can be said that “a convex pentagon does not belong to known Type families” and “a convex pentagon cannot admit tiling” are equivalent. At the stage where these are not truly established [18], it is necessary to confirm whether tiling is truly impossible for a convex pentagon, including whether it is possible to generate only non-periodic tilings. In this study, we assume that Rao’s claim is correct and proceed with the discussion. In the study, to check whether or not the convex pentagon is monotile in Section 2, we do not judge based on only the known Type conditions, and are checking whether tiling is actually possible. On the contrary, if convex pentagons are convex polygonal monotiles that can generate edge-to-edge tilings, it has been proved that they can always generate periodic tilings [11, 12].. Because the set of these polygons is the ASP comprising three types of convex polygons, if two types of polygons are extracted from three types of convex polygons, those two types of polygons cannot admit tilings (but the author has not confirmed this).
@@ 2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into ASP comprising convex polygons
Tile(1,1)11(1,1)( 1 , 1 ), corresponding to $a=b=1$ in Tile$(a,b)$ shown in [9], can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process (see Figure 2). However, as shown in [10], Tile(1,1)11(1,1)( 1 , 1 ) can only generate non-periodic tiling if and only if it does not allow the use of reflected tiles (see Figure 3). Using this property, ASP comprising three types of convex polygons can be obtained from Tile(1,1)11(1,1)( 1 , 1 ). This method divides the interior of Tile(1,1)11(1,1)( 1 , 1 ) into five convex polygons with five or more edges666 Because all triangles and quadrilaterals are monotiles (which do not require the use of reflected tiles to form periodic tilings) the number of edges of convex polygons contained in ASP is at least five. Then, there are no convex polygonal monotiles with seven or more edges [3, 5]., and we have so far identified four methods (Methods 1–4). The tile sets created by the division of the four methods all contain the convex pentagon $P_{1}$, which cannot generate a tiling (that is, no monotile) in the upper-left corner of Figure 4 ($P_{1}$ is convex pentagons with line symmetry and cannot distinguish between the anterior and posterior sides). Methods 1 (Figure 4), 2 (Figure 5), and 4 (Figure 18) divide Tile(1,1)11(1,1)( 1 , 1 ) to create five convex polygons containing three $P_{1}$. In Method 3 (Figure 15), one $P_{1}$ (= $CP_{1}(X_{3})$) is contained in the five convex polygons formed by dividing Tile(1,1)11(1,1)( 1 , 1 ). We note that the tilings generated by the ASP, comprising three types of convex polygons created by Methods 1–4, are non-edge-to-edge.
@@
Figure. 2: Tile(1,1)11(1,1)( 1 , 1 ), and periodic tiling with anterior and posterior sides of Tile(1,1)11(1,1)( 1 , 1 ).
Figure. 3: Non-periodic tiling generated by using only one side of Tile(1,1)11(1,1)( 1 , 1 ).
@@ 2.1 Method 1 and Method 2
Figure. 4: Cases in Method 1.
Figure. 5: Cases in Method 2.
Figure. 6: Periodic tiling with $P_{1}$ and $P_{2}$ with $\alpha=165^{\circ}$ in Method 1 ($\beta=165^{\circ}$ in Method 2).
Figure. 7: Periodic tiling with $P_{1}$ and $P_{3}$ with $\alpha=180^{\circ}$ in Method 1.
@@
Figure. 8: Case where the tile set is not aperiodic in Method 1.
@@
Figure. 9: Concrete example of an ASP that can be created using Method 1 (Part 1).
Figure. 10: Concrete example of an ASP that can be created using Method 1 (Part 2).
Figure. 11: Non-periodic tiling generated by the ASP comprising two types of convex pentagons and one type of convex heptagon in the case of $\alpha=98^{\circ}$ in Method 1.
Figure. 12: Non-periodic tiling generated by the ASP comprising one type of convex pentagon and two types of convex hexagons in the case of $\alpha=124^{\circ}$ in Method 1.
Figure. 13: Combination focusing on vertices with internal angles of $157^{\circ}$ and $150^{\circ}$ in the case of $\alpha=98^{\circ}$ in Method 1.
Figure. 14: Non-periodic tiling generated by the ASP comprising two types of convex pentagons and one type of convex hexagon in the case of $\beta=180^{\circ}$ in Method 2.
@@ 2.2 Method 3
Figure. 15: Case in Method 3.
Figure. 16: Concrete example of an ASP that can be created using Method 3.
Figure. 17: Non-periodic tiling generated by the ASP comprising two types of convex pentagons and one type of convex heptagon in Figure 16.
@@ 2.3 Method 4
Figure. 18: Cases in Method 4.
It can be considered that the tile set of $P_{1}$, $P_{2}$, and $P_{3}$ in the case of $\gamma=75^{\circ}$ in Method 4 (see Figure 19) is not an ASP, because the set contains the convex pentagonal monotile $P_{2}$ belonging to the Type 13 family. However, in this study, we assumed that the reflected tiles cannot be used during all tiling generation processes. $P_{2}$, which belongs to the Type 13 family, requires the use of reflected tiles during the tiling generation process (see Figure 19)888 If convex pentagonal monotiles do not belong to the Types 1 or 3–5 families, then the convex pentagonal monotiles cannot generate tilings without using reflected tiles (see Appendix A). The convex pentagonal monotiles belonging to the Type 13 family cannot belong to the Types 1 or 3–5 families at the same time due to the tile conditions (i.e., the convex pentagonal monotiles belonging to the Type 13 families and those belonging to the Types 1 or 3–5 families are disjoint) [3, 11, 12, 18].. The tile set of the three types of convex pentagons created by this division does not contain reflected $P_{2}$. In other words, tiling cannot be generated with only this $P_{2}$ under the condition that “reflected tiles cannot be used during all tiling generation processes” (we actually confirmed this fact). We examined whether tiling is possible using $P_{1}$, $P_{2}$, and $P_{3}$, as shown in Figure 19, and confirmed that the only combination that can construct a structure in which the tiling does not collapse using $P_{1}$, $P_{2}$, and $P_{3}$ is that of dividing Tile(1,1)11(1,1)( 1 , 1 ) (using the shape of Tile(1,1)11(1,1)( 1 , 1 )), as shown in Figure 19. Therefore, under the condition that the tiles in the tile set cannot be reflected during the tiling generation process, the tile set of the three types of convex pentagons that can be created using the case of $\gamma=75^{\circ}$ in Method 4 is an ASP.
Figure. 19: Case of $\gamma=75^{\circ}$ in Method 4.
Figure. 20: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in in the case of $\gamma=75^{\circ}$ in Method 4.
@@
Figure. 22: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in in the case of $\gamma=180^{\circ}$ in Method 4.
@@ 3 Conclusion
Because the ASPs comprising three types of convex polygons shown in this study are based on Tile(1,1)11(1,1)( 1 , 1 ) shown in [9] and assumes that Tile(1,1)11(1,1)( 1 , 1 ) is the chiral aperiodic monotile shown in [10], it is not confirmed that these sets are actually ASPs. Thus, because [10] is still a preview version, it does not hold if Tile(1,1)11(1,1)( 1 , 1 ) is not a chiral aperiodic monotile (thus far, there is no evidence for such a case).
@@
Currently, there are 15 known families of convex pentagonal monotiles called “Type,” as shown in Figure 23 [2, 3, 4, 6, 7, 11, 12, 18]. The tilings depicted with each convex pentagon in Figure 23 are representative of this Type. A representative tiling can be formed using only the relationships that can be derived from the tile conditions of each Type. As shown in Figure 23, the representative tilings of the Types 2 and 6–15 families use reflected tiles (convex pentagons). In other words, if convex pentagonal monotiles do not belong to the Types 1 or 3–5 families, then the convex pentagonal monotiles cannot generate tilings without using reflected tiles.
However, Type families of convex pentagonal monotiles are not necessarily “disjoint.” That is, there are some convex pentagonal monotiles that belong to more than one Type family [11, 13, 14, 15]. For example, the convex pentagonal monotile in Figure 24 belongs to both the Type 1 and Type 7 families [13, 14]. The convex pentagonal monotile belonging only to the Type 7 family can form a representative tiling of the Type 7 family, as shown in Figure 23, and requires reflected tiles to form the tiling. By contrast, the convex pentagonal monotiles in Figure 24, which belong to both the Type 1 and Type 7 families, can form the tiling of the Type 7 family, but they can also form the tiling of the Type 1 family without using reflected tiles. In other words, the convex pentagonal monotiles belonging to both the Type 1 and Type 7 families in Figure 24 can generate tiling without using reflected tiles. It should be noted that some convex pentagonal monotiles can form tilings other than the representative tilings of each Type [13, 14, 15].
@@
Figure. 23: Fifteen Type families of convex pentagonal monotiles. Each of the convex pentagonal monotiles is defined by some conditions between the lengths of the edges and the magnitudes of the angles, but some degrees of freedom remain. For example, a convex pentagonal monotile belonging to Type 1 family satisfies that the sum of three consecutive angles is equal to $360^{\circ}$. This condition for Type 1 family is expressed as $A+B+C=360^{\circ}$ in this figure. The pentagonal monotiles of Types 14 and 15 families have one degree of freedom, that of size. For example, the value of $C$ of the pentagonal monotile of Type 14 family is $\cos^{-1}((3\sqrt{57}-17)/16)\approx 1.2099\;\mbox{rad}\approx 69.32^{\circ}$. The gray region in each tiling indicates a translation unit (a unit that can generate a periodic tiling by translation only).
Figure. 24: Convex pentagonal monotile belonging to both the Type 1 and Type 7 families, the tilings of Type 1 family and the tiling of Type 7 with the monotiles.
Figure. 25: Periodic tiling with convex pentagonal monotile belonging to both the Type 2 and Type 4 families.
@@
Three families of convex hexagonal monotiles are known, called “Type,” as shown in Figure 26 [2, 3, 4, 8, 12]. The tilings depicted with each convex hexagon in Figure 26 are representative tilings of Type. Representative tiling can be formed using only those relations that can be derived from the tile conditions of each Type. As shown in Figure 26, the representative tiling of the Type 2 family uses reflected tiles (convex hexagons). If convex hexagonal monotiles do not belong to the Type 1 or Type 3 families, then the convex hexagonal monotiles cannot generate tilings without using reflected tiles. The Type families of convex hexagonal monotiles are not necessarily “disjoint.” For example, a regular hexagon belongs to all Type families.
@@
Figure. 26: Three Type families of convex hexagonal monotiles. The gray region in each tiling indicates a translation unit (a unit that can generate a periodic tiling by translation only).
Figure. 27: Periodic tilings with convex hexagonal monotiles belonging to the Type 1 family, created by the division of Method 1.
@@
Figure. 28: Periodic tiling with $P_{1}$, $P_{2}$, and reflected $P_{2}$ in the case of $\alpha=180^{\circ}$ in Method 1 (see Figure 4(g)). The reflected $P_{2}$ are given an asterisk mark “*.”
Appendix Appendix D ASPs comprising four or five types of convex polygons999 The contents (the results of study) in this appendix were obtained prior to the main results of this manuscript.

### v5

@@ 1 Introduction
A tiling (or tessellation) of the plane is a collection of sets, called tiles, that covers the plane without gaps or overlaps, except for the boundaries of the tiles. The term “tile” refers to a topological disk, whose boundary is a simple closed curve.
A tiling exhibits periodicity if its translation by a non-zero vector coincides with itself; a tiling is considered periodic if it coincides with its translation by two linearly independent vectors. However, in this study, a tiling with periodicity is referred to as periodic, and a tiling without periodicity is referred to as non-periodic.
In general, when considering tilings, there are only a finite number of tile shapes that can be used. These finite-shape diagrams are called prototiles, and a set of prototiles admits tilings of the plane. The tiling generated by a set of prototiles can cover the plane infinitely. A set of prototiles is said to be aperiodic if copies of the prototiles can be assembled into tilings of the plane such that all tilings with the prototiles are non-periodic [2, 4, 22].
The most well-known problem regarding a set of prototiles (set of tiles, , tile set) that are aperiodic, that is, “aperiodic set of prototiles (aperiodic set of tiles)” (hereafter abbreviated as “ASP”), is “Is there a single aperiodic prototile (with or without a matching condition111 Matching conditions specify how tiles must connect to form a valid tiling, which can sometimes be represented by assigning colors or orientations to specific edges of the prototiles. ), that is, one that admits only non-periodic tilings by congruent copies?” [2]. Smith et al. presented solutions to this problem for concave diagram tiles without matching conditions [12, 13]. In [2], another major problem concerning ASP was presented as follows:
@@
Figure 1: Tiling (d) generated by an aperiodic set of prototiles created by Ammann, which used three convex polygons with no matching condition on the edges. As shown in (c), this is a recomposition by Penrose tiles (rhombuses) and is obtained by markings of using the points $J$, $K$, and $L$ as shown in (a) and (b). These markings are completely determined by the choice of the point $J$, as we must have GL = DJ, FL = CJ =EK, and HK = BJ. For aperiodicity, the point $J$ must be chosen so that DJ, CJ, BJ, and KL are of different lengths [4, 16].
The “ASP comprising three convex polygons (= the set of three convex polygons that are aperiodic with no matching condition on the edges)” in the above problem is known for the two convex pentagons and one convex hexagon created by Ammann, as shown in Figure 1, based on the Penrose tiles222 The Penrose tiles are considered to be an ASP comprising two convex polygons (rhombuses) with matching condition on the edges [2, 4, 16]. [4, 16]. In the tiling of the two convex pentagons and one convex hexagon shown in Figure 1, the reflected (flipped) convex pentagons and hexagon were not used. The tiling is edge-to-edge333 A tiling by convex polygons is edge-to-edge if any two convex polygons in a tiling are either disjoint or share one vertex or an entire edge in common. Then other case is non-edge-to-edge [4, 15, 16]..
@@
According to Smith et al. [12], Tile$(a,b)$, which can generate only non-periodic tiling, is an “aperiodic monotile,” in which all tiles are of the same size and shape; however, the tiling generation process requires the use of reflected tiles, and both anterior and posterior side tiles must be treated as the same type. However, according to Smith et al. [13], the tiles “Spectres,” which can generate only non-periodic tiling, do not require the use of reflected tiles during the tiling generation process444 Spectres can generate only non-periodic tiling using only one side, either anterior or posterior, even if it allows the use of reflected tiles during the tiling generation process.. The tiles are referred to as “chiral aperiodic monotiles.”
@@
Following this policy, the case in [12] is “ASP comprising two types of concave polygons,” and the case in [13] is “ASP comprising one type of concave diagram555 Tile(1,1)11(1,1)( 1 , 1 ) corresponding to $a=b=1$ of Tile$(a,b)$ shown in [12] is a concave polygon and a concave diagram, and Spectres in [13] are concave diagrams. From a viewpoint that does not distinguish between anterior and posterior sides, both cases of [12] and [13] are considered “ASP comprising one type of concave diagram.” Note that Tile(1,1)11(1,1)( 1 , 1 ) with 13-edges is considered an equilateral 14-edges polygon.”. Moreover, the case shown by Ammann in Figure 1 is “ASP comprising three types of convex polygons (two types of convex pentagons and one type of convex hexagon).”
@@
The two types of convex pentagons and one type of convex hexagon shown in Figure 1 can be identified as not belonging to known Type families of convex pentagonal monotiles (see Appendix A) or convex hexagonal monotiles (see Appendix B), respectively666 If Rao’s claim in [9] is formally established as correct, there will be no ASP for convex polygonal monotiles (i.e., there does not exist ASP comprising one type of convex polygon with no matching condition on the edges), and furthermore, it can be said that “a convex pentagon does not belong to known Type families” and “a convex pentagon cannot admit tiling” are equivalent. At the stage where these are not truly established [24], it is necessary to confirm whether tiling is truly impossible for a convex pentagon, including whether it is possible to generate only non-periodic tilings. In this study, we assume that Rao’s claim is correct and proceed with the discussion. However, in Section 2 (and Appendix D) of this study, to check whether or not a convex polygon is a monotile, we do not judge based on only the tile conditions of known Types, and are checking whether tiling is actually possible. Note that if convex polygons are convex polygonal monotiles that can generate edge-to-edge tilings, it has been proved that they can always generate periodic tilings [15, 16]. . Because the set of these polygons is the ASP comprising three types of convex polygons, if two types of polygons are extracted from three types of convex polygons, those two types of polygons cannot admit tilings (but the author has not confirmed this).
@@ 2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into ASP comprising convex polygons
Tile(1,1)11(1,1)( 1 , 1 ), corresponding to $a=b=1$ in Tile$(a,b)$ shown in [12], can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process (see Figure 2). However, as shown in [13], Tile(1,1)11(1,1)( 1 , 1 ) can only generate non-periodic tiling if and only if it does not allow the use of reflected tiles (see Figure 3). Using this property, ASP comprising three types of convex polygons can be obtained from Tile(1,1)11(1,1)( 1 , 1 ). This method divides the interior of Tile(1,1)11(1,1)( 1 , 1 ) into five convex polygons with five or more edges777 Because all triangles and quadrilaterals are monotiles (which do not require the use of reflected tiles to form periodic tilings) the number of edges of convex polygons contained in ASP is at least five. Then, there are no convex polygonal monotiles with seven or more edges [3, 4, 7, 5, 10, 25]. , and we have so far identified four methods (Methods 1–4). The tile sets created by the division of the four methods all contain the convex pentagon $P_{1}$, which cannot generate a tiling (that is, no monotile) in the upper-left corner of Figure 4 ($P_{1}$ is convex pentagons with line symmetry and cannot distinguish between the anterior and posterior sides). Methods 1 (Figure 4), 2 (Figure 5), and 4 (Figure 18) divide Tile(1,1)11(1,1)( 1 , 1 ) to create five convex polygons containing three $P_{1}$. In Method 3 (Figure 15), one $P_{1}$ (= $CP_{1}(X_{3})$) is contained in the five convex polygons formed by dividing Tile(1,1)11(1,1)( 1 , 1 ). We note that the tilings generated by the ASP, comprising three types of convex polygons created by Methods 1–4, are non-edge-to-edge.
@@
Figure 2: Tile(1,1)11(1,1)( 1 , 1 ), and periodic tiling with anterior and posterior sides of Tile(1,1)11(1,1)( 1 , 1 ).
Figure 3: Non-periodic tiling generated by using only one side of Tile(1,1)11(1,1)( 1 , 1 ).
@@ 2.1 Method 1 and Method 2
Figure 4: Cases in Method 1.
Figure 5: Cases in Method 2.
Figure 6: Periodic tiling with $P_{1}$ and $P_{2}$ with $\alpha=165^{\circ}$ in Method 1 ($\beta=165^{\circ}$ in Method 2).
Figure 7: Periodic tiling with $P_{1}$ and $P_{3}$ with $\alpha=180^{\circ}$ in Method 1.
@@
Figure 8: Case where the tile set is not aperiodic in Method 1.
@@
Figure 9: Concrete example of an ASP that can be created using Method 1 (Part 1).
Figure 10: Concrete example of an ASP that can be created using Method 1 (Part 2).
Figure 11: Non-periodic tiling generated by the ASP comprising two types of convex pentagons and one type of convex heptagon in the case of $\alpha=98^{\circ}$ in Method 1.
Figure 12: Non-periodic tiling generated by the ASP comprising one type of convex pentagon and two types of convex hexagons in the case of $\alpha=124^{\circ}$ in Method 1.
Figure 13: Combination focusing on vertices with internal angles of $157^{\circ}$ and $150^{\circ}$ in the case of $\alpha=98^{\circ}$ in Method 1.
Figure 14: Non-periodic tiling generated by the ASP comprising two types of convex pentagons and one type of convex hexagon in the case of $\beta=180^{\circ}$ in Method 2.
@@ 2.2 Method 3
Figure 15: Case in Method 3.
Figure 16: Concrete example of an ASP that can be created using Method 3.
Figure 17: Non-periodic tiling generated by the ASP comprising two types of convex pentagons and one type of convex heptagon in Figure 16.
@@ 2.3 Method 4
Figure 18: Cases in Method 4.
It can be considered that the tile set of $P_{1}$, $P_{2}$, and $P_{3}$ in the case of $\gamma=75^{\circ}$ in Method 4 (see Figure 19) is not an ASP, because the set contains the convex pentagonal monotile $P_{2}$ belonging to the Type 13 family. However, in this study, we assumed that the reflected tiles cannot be used during all tiling generation processes. $P_{2}$, which belongs to the Type 13 family, can form the (representative) tiling of the Type 13 family shown in Figure 19 by using reflected $P_{2}$. However, the tile set of the three types of convex pentagons created by this division does not contain the reflected $P_{2}$. We actually confirmed that tiling cannot be generated with only this $P_{2}$ under the condition that “reflected tiles cannot be used during all tiling generation processes”999 If a convex pentagonal monotile belongs to at least one of the Type 1 or 3–6 families, it can generate tilings without the use of reflected tiles (see Appendix A). The convex pentagonal monotiles belonging to the Type 13 families and those belonging to the Type 1 or 3–6 families are disjoint [4, 15, 16, 17, 20, 24]. . We examined whether tiling is possible using $P_{1}$, $P_{2}$, and $P_{3}$, as shown in Figure 19, and confirmed that the only combination that can construct a structure in which the tiling does not collapse using $P_{1}$, $P_{2}$, and $P_{3}$ is that of dividing Tile(1,1)11(1,1)( 1 , 1 ) (using the shape of Tile(1,1)11(1,1)( 1 , 1 )), as shown in Figure 19. Therefore, under the condition that the tiles in the tile set cannot be reflected during the tiling generation process, the tile set of the three types of convex pentagons that can be created using the case of $\gamma=75^{\circ}$ in Method 4 is an ASP.
@@
Figure 19: Case of $\gamma=75^{\circ}$ in Method 4.
Figure 20: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in in the case of $\gamma=75^{\circ}$ in Method 4.
@@
Figure 22: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in in the case of $\gamma=180^{\circ}$ in Method 4.
@@ 3 Conclusion
Because the ASPs comprising three types of convex polygons shown in this study are based on Tile(1,1)11(1,1)( 1 , 1 ) shown in [12] and assumes that Tile(1,1)11(1,1)( 1 , 1 ) is the chiral aperiodic monotile shown in [13], it is not confirmed that these sets are actually ASPs.
@@
Currently, there are 15 known families of convex pentagonal monotiles, each labeled as a “Type,” as shown in Figure 23 [3, 4, 6, 8, 9, 11, 14, 15, 16, 17, 20, 24, 25]. The tilings depicted with each convex pentagon in Figure 23 are representative of each Type family. A representative tiling can be formed using only the relationships that can be derived from the tile conditions of each Type. As shown in Figure 23, the representative tilings of the Types 2 and 7–15 families use reflected tiles (convex pentagons). By contrast, as shown in Figure 23, the representative tilings of the Types 1 and 3–6 families do not use reflected tiles. Thus, if a convex pentagonal monotile belongs to at least one of the Type 1 or 3–6 families, it can generate tilings without the use of reflected tiles [20].
Figure 23: Fifteen Type families of convex pentagonal monotiles. Each of the convex pentagonal monotiles is defined by some conditions between the lengths of the edges and the magnitudes of the angles, but some degrees of freedom remain. For example, convex pentagonal monotiles belonging to the Type 1 family (also referred to as convex pentagonal monotiles belonging to Type 1, or simply as convex pentagonal monotiles of Type 1) satisfy the property the sum of three consecutive angles is equal to $360^{\circ}$. In this figure, this property is expressed as $A+B+C=360^{\circ}$, which represents the tile conditions of Type 1 (the Type 1 family). The pentagonal monotiles of Types 14 and 15 have one degree of freedom, that of size. For example, the value of angle $C$ in the pentagonal monotile of Type 14 is $\cos^{-1}\bigl{(}\bigl{(}3\sqrt{57}-17\big{)}/16\big{)}\approx 1.2099\;\mbox{rad}\approx 69.32^{\circ}$. The gray region in each tiling indicates a translation unit (a unit that can generate a periodic tiling through translation alone).
The Type families of convex pentagonal monotiles are not necessarily “disjoint.” That is, there are some convex pentagonal monotiles that belong to more than one Type family [15, 18, 19, 20, 21]. For example, the convex pentagonal monotile in Figure 24 belongs to both the Type 1 and Type 7 families [18, 19]. If $A=86^{\circ}$ for the tile conditions of Type 7 in Figure 23, the convex pentagonal monotile with $A=86^{\circ}$ belongs only to the Type 7 family and can only form the (representative) tiling of the Type 7 family, as depicted in Figure 23. By contrast, the convex pentagonal monotiles in Figure 24, which belong to both the Type 1 and Type 7 families, can form the tiling of the Type 7 family, but they can also form the tiling of the Type 1 family without using reflected tiles. In other words, the convex pentagonal monotiles belonging to both the Type 1 and Type 7 families in Figure 24 can generate tiling without using reflected tiles. It should be noted that some convex pentagonal monotiles can form tilings other than the representative tilings of each Type [18, 19, 20, 21].
@@
Figure 24: Convex pentagonal monotile belonging to both the Type 1 and Type 7 families, the tilings of the Type 1 family and the tiling of the Type 7 family with the convex pentagon.
Figure 25: Periodic tiling with convex pentagonal monotile belonging to both the Type 2 and Type 4 families.
@@
Three families of convex hexagonal monotiles are known, called “Type,” as shown in Figure 26 [1, 3, 4, 6, 10, 16, 25]. The tilings depicted with each convex hexagon in Figure 26 are representative tilings of each Type family. Representative tiling can be formed using only those relations that can be derived from the tile conditions of each Type. As shown in Figure 26, the representative tiling of the Type 2 family uses reflected tiles (convex hexagons). If convex hexagonal monotiles do not belong to the Type 1 or Type 3 families, then the convex hexagonal monotiles cannot generate tilings without using reflected tiles. The Type families of convex hexagonal monotiles are not “disjoint.” For example, a regular hexagon belongs to all Type families.
@@
Figure 26: Three Type families of convex hexagonal monotiles. The gray region in each tiling indicates a translation unit (a unit that can generate a periodic tiling by translation only).
Figure 27: Periodic tilings with convex hexagonal monotiles belonging to the Type 1 family, created by the division of Method 1.
@@
Figure 28: Periodic tiling with $P_{1}$, $P_{2}$, and reflected $P_{2}$ in the case of $\alpha=180^{\circ}$ in Method 1 (see Figure 4(g)). The reflected $P_{2}$ are given an asterisk mark “*.”
Appendix Appendix D ASPs comprising four or five types of convex polygons101010 The contents (the results of study) in this appendix were obtained prior to the main results of this manuscript.